A triangle has sides of squareroot of 2 and 3. Which could not be the length of the third side if it is a right triangle?


squareroot of 7 OR squareroot of 11 OR squareroot of 13.

Answer by stanbon(75887)

A triangle has sides of squareroot of 2 and 3. Which could not be the length of the third side if it is a right triangle?
---
Let the 3rd side be "x":
If x is the hypotenuse, x^2 = (sqrt(2)^2 + (3)^2
x^2 = 2 + 9
x = sqrt(11)
---------------------------------
If x is one of the legs:
3^2 = x^2 + (sqrt(2))^2
9 = x^2 + 2
x^2 = 7
x = sqrt(7)
---------------------------------
Cheers,
Stan H.
=================================

squareroot of 7 OR squareroot of 11 OR squareroot of 13.